You want to know how many years it will take you to accumulate $1,216.65 if you invest $1,000.00 in an investment that pays annual interest at 4%.
a) Use the NPER function in Excel to calculate the number of periods if you invested $1,000 today in an investment that pays 4% and will accumulate $1,216.
b) Now use the present value or future value tables as shown in the text to calculate the number of years is will take.
Answer #1
Present value ($) 1000
Future value ($) 1216.65
Rate of interest 4%
Nper function =nper(4%,0,1000,-1216.65,0)
Number of years to reach the future value 5.00
Using present and future value tables
Let number of years = x
Present value ($) 1000
Future value ($) 1216.65
Rate of interest 4%
We know that 1000*Future value factor at 4% for x years = 1216.65
Future value factor at 4% for x years = 1216.65/1000 = 1.2166
Seeing the future value table, it can be seen that 1.2166 appears at 4% in 5 years. Hence, the time period is 5 years